### v1

# Dynamic Contrastive Knowledge Distillation for Efficient Image Restoration

Abstract
Knowledge distillation (KD) is a valuable yet challenging approach that enhances a compact student network by learning from a high-performance but cumbersome teacher model. However, previous KD methods for image restoration overlook the state of the student during the distillation, adopting a fixed solution space that limits the capability of KD. Additionally, relying solely on L1-type loss struggles to leverage the distribution information of images. In this work, we propose a novel dynamic contrastive knowledge distillation (DCKD) framework for image restoration. Specifically, we introduce dynamic contrastive regularization to perceive the student’s learning state and dynamically adjust the distilled solution space using contrastive learning. Additionally, we also propose a distribution mapping module to extract and align the pixel-level category distribution of the teacher and student models. Note that the proposed DCKD is a structure-agnostic distillation framework, which can adapt to different backbones and can be combined with methods that optimize upper-bound constraints to further enhance model performance. Extensive experiments demonstrate that DCKD significantly outperforms the state-of-the-art KD methods across various image restoration tasks and backbones.
Introduction
Image restoration aims to recover high-quality images from low-quality ones degraded by processes such as sub-sampling, blurring, and rain streaks. It is a highly challenging ill-posed inverse problem since crucial content information is missing during degradation. Recently, convolutional neural networks (CNNs) (Dong et al. 2015; Lim et al. 2017; Zhang et al. 2018) and Transformers (Liang et al. 2021; Chen et al. 2021; Wang et al. 2022; Zamir et al. 2022) have been extensively investigated for designing various models, achieving remarkable success in image restoration. However, these models demand high resources and exhibit inefficiencies, making deployment on resource-constrained devices challenging (Zhang et al. 2021). To facilitate their real-world applications, there is a growing research focus on compressing image restoration models.
Figure 1: Difference between our DCKD and existing KD methods. (a) The Vanilla KD method overlooks the information from negative images as a lower bound. (b) Existing contrastive KD methods adopt a fixed lower bound that limits the capability of KD. (c) Our DCKD introduces dynamic contrastive regularization to perceive the student’s learning state and dynamically adjust the distilled solution space.
Knowledge distillation (KD) is an effective model compression method that transfers knowledge from a cumbersome teacher model to a lightweight student model. This process allows the student models to inherit the capabilities of the teacher models, resulting in significant performance improvements while reducing computational and storage requirements. KD has gained broad recognition for its excellent performance and broad applicability. It also can be combined with other model compression techniques, such as quantization (Du et al. 2021; Ayazoglu 2021; Hong et al. 2022), pruning (Fan et al. 2020; Wang et al. 2021a; Oh et al. 2022), compact architecture design (Ahn, Kang, and Sohn 2018; Zhang et al. 2022; Chen et al. 2022a), and neural architecture search (NAS) (Gou et al. 2020; Kim et al. 2021; He et al. 2022), to enhance the compactness of student models further.
Since KD has been well-established in natural language processing (Hahn and Choi 2019; Sanh et al. 2019; Fu et al. 2021) and high-level vision tasks (Touvron et al. 2021; Lin et al. 2022; Chen et al. 2022b), researchers have been investigating KD for image restoration methods (He et al. 2020; Lee et al. 2020; Wang et al. 2021c; Zhang et al. 2023; Li et al. 2024; Jiang et al. 2024; Zhang et al. 2024). However, these methods adopt a fixed solution space, which limits their adaptability to the evolving state of the student model during the distillation process. As illustrated in Fig. 1 (a), the vanilla distillation method (Hinton, Vinyals, and Dean 2015) only constrains the upper bound of the solution space. Although existing works (Zhang et al. 2023; Jiang et al. 2024) explore more effective and diverse upper bounds, the lack of constraints on the lower bound of the output image increases the difficulty of optimizing the solution space. This often generates low-quality images with artifacts, color distortion, and blurring. As illustrated in Fig. 1 (b), CSD (Wang et al. 2021c) introduces contrastive learning to design lower-bound constraints, significantly enhancing the transfer of knowledge from the teacher. However, in the later stages of training, the student anchor moves far from the lower bound, leading to a diminished constraint effect.
To address this problem, we propose a novel dynamic contrastive knowledge distillation framework named DCKD. Specifically, we first propose the dynamic contrastive regularization, which generates dynamic lower bound constraints. In addition, we also propose a distribution mapping module (DMM) to extract and align the pixel-level category distribution between the output of the teacher and student. Compared with previous image restoration distillation methods that primarily relied on L1 loss, DMM successfully introduces category distribution information distillation into low-level vision tasks. DCKD not only adapts to various backbones but also can be combined with methods (Zhang et al. 2023; Li et al. 2024; Jiang et al. 2024) that improve the upper bound of the solution space to further enhance distillation performance. We validate the effectiveness of DCKD across multiple image restoration tasks, including image super-resolution, deblurring, and deraining.
Overall, our main contributions are as follows:
  • We propose a dynamic contrastive distillation framework (DCKD), which can perceive the student’s learning state and dynamically optimize the lower bound of the solution space.
  • We introduce a distribution mapping module to leverage category distribution information distillation, which has been significantly ignored in previous KD works for image restoration.
  • Extensive experiments across various image restoration tasks demonstrate that the proposed DCKD framework significantly outperforms previous methods.
Related Work
Image Restoration
Since the pioneering works SRCNN (Dong et al. 2015) and DnCNN (Zhang et al. 2017) are firstly to employ CNNs for image restoration, various works (Lim et al. 2017; Nah, Hyun Kim, and Mu Lee 2017a; Lefkimmiatis 2017; Li et al. 2018; Ren et al. 2019; Chen et al. 2022a) have been proposed to improve the performance by increasing the parameters. Recently, Transformer-based methods (Chen et al. 2021; Liang et al. 2021; Chen et al. 2022c, 2023) have leveraged self-attention mechanisms to capture long-range dependencies, leading to significant performance improvements in image restoration. To reduce computational overhead, SAFMN (Sun et al. 2023) enhances model efficiency by utilizing spatially adaptive feature pyramid attention maps. Restormer (Zamir et al. 2022) designs channel self-attention, which is more efficient than spatial self-attention. Although lightweight designs of CNN and Transformer architectures significantly reduce computational overhead, they still face challenges regarding direct deployment on resource-constrained platforms.
Knowledge Distillation for Image Restoration
Knowledge distillation aims to significantly reduce deployment costs while improving the performance of student models by emulating the behavior of teacher models (Hinton, Vinyals, and Dean 2015; Gou et al. 2021; Lee et al. 2020). In recent years, numerous works have focused on knowledge distillation for image restoration. He et al. proposed FAKD to align the spatial affinity matrix of the feature maps between the teacher and student models (He et al. 2020).To alleviate the semantic differences between features of the teacher and student, Li et al. proposed MiPKD, which achieves feature and stochastic network block mixture in latent space (Li et al. 2024). Jiang et al. proposed MTKD that designs a composite output from multiple teachers to provide the student with a more robust teacher model (Jiang et al. 2024). However, these KD methods primarily focus on improving the upper bound of the model’s solution space, without leveraging the distribution information of the images.
Contrastive Learning for Knowledge Distillation
Recently, many researchers have explored the combination of contrastive learning and knowledge distillation to construct a comprehensive solution space (Tian, Krishnan, and Isola 2019; Xu et al. 2020; Yang et al. 2023). Wang et al. introduced contrastive learning in knowledge distillation, utilizing other images within the same dataset to construct a lower bound for the solution space (Wang et al. 2021b). Similarly, CSD (Wang et al. 2021c) proposed a contrastive self-distillation method, utilizing different images within the batch to provide lower bound constraints. Luo et al. enriched the lower bound constraints in the image deraining task by altering the direction of falling raindrops for the student model (Luo et al. 2023). However, these methods employ a fixed solution space, which leads to a weakening of the lower-bound constraint when the student anchor moves away from the lower bound in the later stages of training. Different from these methods, DCKD proposes a dynamic lower-bound constraint that progressively narrows the solution space, and can also be combined with enhanced upper-bound approaches. Additionally, DMM is proposed to extract and align the pixel-level category distribution information between the teacher and student networks.
Figure 2: Illustration of the proposed Dynamic Contrastive Knowledge Distillation framework. Our DCKD consists of two parts: (a) Dynamic Contrastive Regularization (DCR), and (b) Distribution Mapping Module (DMM).
Methodology
Preliminaries
Given a low-quality image $I_{LQ}$ as input, the image restoration (IR) model $\mathcal{F}(\cdot)$ generates the corresponding high-quality image $I_{HQ}$, which can be formulated as:
$I_{HQ}=\mathcal{F}(I_{LQ};\theta),$ (1)
where $\theta$ represents the model parameters. The IR model $\mathcal{F}(\cdot;\theta)$ is typically optimized using the L1 norm reconstruction loss, which is defined as:
$\mathcal{L}_{rec}=\|I_{HQ}-I_{GT}\|_{1},$ (2)
where the $I_{GT}$ is the ground-truth image. The vanilla knowledge distillation method adds a KD loss to minimize the difference between the student model and the teacher model:
$\mathcal{L}_{kd}=\|\mathcal{F}_{S}(I_{LQ};\theta_{s})-\mathcal{F}_{T}(I_{LQ};\theta_{t})\|_{1},$ (3)
where $\mathcal{F}_{S}(I_{LQ};\theta_{s})$ and $\mathcal{F}_{T}(I_{LQ};\theta_{t})$ represent the outputs of student and teacher models, respectively.
Dynamic Contrastive Regularization
Previous KD methods (Hinton, Vinyals, and Dean 2015; Wang et al. 2021c) employed a fixed solution space, which causes the lower-bound constraints on the student model to weaken in the later stages of training. To address this problem, we propose dynamic contrastive regularization (DCR), which dynamically adjusts the solution space based on the student model’s state.
As shown in Fig. 2 (a), we first feed the input image $I_{LQ}$ into the dynamic negative sample generator, which consists of the degrading module and the history model. The degrading module applies random degrading operations on $I_{LQ}$, generating $N$ different degraded images $I_{dirty}$. The history model $\mathcal{F}_{S}^{his}$ then reconstructs these degraded images based on the historical state of the student model, producing $N$ different negative images $I_{Neg}$ as the lower-bound of the solution space:
$I_{Neg}^{1},\cdots,I_{Neg}^{N}=\mathcal{F}_{S}^{his}(\mathbb{D}_{1}(I_{LQ}),\cdots,\mathbb{D}_{N}(I_{LQ})),$ (4)
where $\mathbb{D}_{N}(\cdot)$ and $I_{Neg}^{N}$ represent the $N$-th type of random degradation and the corresponding negative image, respectively.
For the upper-bound of the solution space, we not only rely on ground-truth $I_{GT}$ to optimize the student output, as described in Eq. 2 but also consider the output of teacher model $I_{HQ}^{T}$ as the positive image $I_{Pos}$. At this point, we have defined the upper-bound and the dynamic lower-bound of the model, allowing us to construct a novel dynamic contrastive loss. We employ the pre-trained VQGAN (Esser, Rombach, and Ommer 2021) as the feature encoder. The dynamic contrastive loss is formulated as follows:
$\mathcal{L}_{dcl}=\sum_{i=1}^{L}\lambda_{i}\frac{||f_{i}^{Anc}-f_{i}^{Pos}||_{1}}{\sum_{j=1}^{N}||f_{i}^{Anc}-f_{i,j}^{Neg}||_{1}},$ (5)
where $f^{Anc}$, $f^{Pos}$, and $f^{Neg}$ represent the features extracted from the output of the student model, positive image and negative images, respectively. $i$ denotes the $i$-th layer of Encoder. $\lambda_{i}$ is the balancing weight for $i$-th layer.
To better capture the state of the student model, we introduce exponential moving averages (EMA) to update the history model:
$\theta_{his}=\alpha\theta_{his}+(1-\alpha)\theta_{stu},~{}~{}\mathrm{s.t.}~{}~{}t\%s=0,$ (6)
where $\theta_{his}$ and $\theta_{stu}$ represent the parameters of the history model and the current student model, respectively. $\alpha$ is the attenuation rate, $t$ is the current iteration, and $s$ denotes the update step. During training, the update step $s$ gradually increases.
Distribution Mapping Module
Existing image restoration KD methods (Li et al. 2024; Jiang et al. 2024) only rely on L1-type loss to align the teacher and student models, thereby overlooking the distribution information of image content. However, high-level task KD methods (Hinton, Vinyals, and Dean 2015; Park et al. 2019; Huang et al. 2022) that align the entire output distribution of the teacher and student networks fail in low-level vision tasks. To address this, we design a distribution mapping module (DMM) to extract and align pixel-wise image distribution information, which is well-suited for pixel-level image restoration tasks.
Figure 3: Illustration of the Image-level Distribution and our Pixel-level Distribution (DMM).
As shown in Fig. 2 (b), we employ a pre-trained image encoder to extract deep features $F^{T}$ and $F^{S}$ from the output images of the teacher model $I_{HQ}^{T}$ and the student model $I_{HQ}^{S}$, respectively:
$F^{T}=\operatorname{Encoder}\left(I_{HQ}^{T}\right),F^{S}=\operatorname{Encoder}\left(I_{HQ}^{S}\right).$ (7)
We assume that high-level KD methods struggle to provide fine-grained distribution constraints, which are crucial for restoring image details. Inspired by VQGAN (Esser, Rombach, and Ommer 2021), we use the codebook $e$ pre-trained on ImageNet (Krizhevsky, Sutskever, and Hinton 2012) to obtain the pixel-wised category distribution as illustrated in Fig. 3. We also employ the corresponding VQGAN as the image encoder. The codebook further transforms the extracted deep features $F^{T}$ and $F^{S}$ into category distributions can be formulated as:
$C^{T}=\psi(\|F^{T}-e\|_{2}^{2}),C^{S}=\psi(\|F^{S}-e\|_{2}^{2}),$ (8)
where $C^{T}$ and $C^{S}$ represent the pixel-wised category distributions of the teacher and student models, respectively. $\psi$ denotes the softmax operation.
Finally, we employ cross-entropy loss to align $C^{T}$ and $C^{S}$:
$\mathcal{L}_{ce}=-\sum_{i=1}^{M}C_{i}^{T}\log{C_{i}^{S}},$ (9)
where $C_{i},i=1,2,\cdots M$ is category i. $M$ is the total number of categories.
Overall Loss
Following (Hinton, Vinyals, and Dean 2015; Zhang et al. 2023; Li et al. 2024), our DCKD also compute the reconstruction loss $\mathcal{L}_{res}$ in Eq. 2 and the vanilla distillation loss $\mathcal{L}_{kd}$ in Eq. 3. In addition, the dynamic contrastive loss $\mathcal{L}_{dcl}$ in Eq. 5 and the cross-entropy loss $\mathcal{L}_{ce}$ in Eq. 9 are also accumulated. The overall loss function can be expressed as:
$\mathcal{L}=\mathcal{L}_{rec}+\mathcal{L}_{kd}+\lambda_{dcl}\mathcal{L}_{dcl}+\lambda_{ce}\mathcal{L}_{ce},$ (10)
where the $\lambda_{dcl}$ and $\lambda_{ce}$ are the balancing parameters. The teacher model and the encoder are frozen during the training stage.
Experiments
Experimental Settings
Teacher Backbones
The proposed DCKD is evaluated on three image restoration tasks, image super-resolution, image deblurring, and image draining. Following (Li et al. 2024), we verify the effectiveness of DCKD on Transformer-based SwinIR (Liang et al. 2021) and CNN-based RCAN (Zhang et al. 2018) in image super-resolution. For image deblurring, we use NAFNet (Chen et al. 2022a) and Restormer (Zamir et al. 2022) as the teacher backbones. For image deraining, we employ Restormer (Zamir et al. 2022) as the backbone. The configuration details of teacher and student models are presented in the Appendix.
Scale Method Set5 Set14 BSD100 Urban100
PSNR PSNR PSNR PSNR
$\times$2 Teacher 38.36 34.14 32.45 33.40/
Scratch 38.00 33.56 32.19 32.05
Logits 38.04 33.61 32.22 32.09
FAKD 38.03 33.63 32.21 32.06
MiPKD 38.14 33.76 32.29 32.46
DCKD 38.17 33.85 32.31 32.59
$\times$3 Teacher 34.89 30.77 29.37 29.29
Scratch 34.41 30.43 29.12 28.20
Logits 34.44 30.45 29.14 28.23
FAKD 34.42 30.42 29.12 28.18
MiPKD 34.53 30.52 29.19 28.47
DCKD 34.59 30.57 29.22 28.63
$\times$4 Teacher 32.72 28.94 27.83 27.07
Scratch 32.31 28.67 27.61 26.15
Logits 32.27 28.67 27.62 26.15
FAKD 32.22 28.65 27.61 26.09
MiPKD 32.39 28.76 27.68 26.37
DCKD 32.49 28.82 27.72 26.53
$\mathrm{\textbf{DCKD}}^{*}$ 32.51 28.88 27.74 26.62
Table 1: Quantitative comparison of the benchmark datasets for the SwinIR model. The best and second-best performances are highlighted in bold and underlined, respectively. DCKD* adds the optimized upper bound constraint.
Datasets and Evaluation
For image super-resolution, DCKD is trained using 800 images from the DIV2K dataset (Timofte et al. 2017) and evaluated on four benchmark datasets: Set5 (Bevilacqua et al. 2012), Set14 (Zeyde, Elad, and Protter 2012), BSD100 (Martin et al. 2001), and Urban100 (Huang, Singh, and Ahuja 2015). For image deblurring, the models are trained and tested both on the GoPro dataset (Nah, Hyun Kim, and Mu Lee 2017b). For image deraining, we train DCKD on 13,712 clean-rainy image pairs collected from multiple datasets (Fu et al. 2017; Yang et al. 2017; Zhang, Sindagi, and Patel 2019; Li et al. 2016) and evaluate it on Test100 (Zhang, Sindagi, and Patel 2019), Rain100H (Yang et al. 2017), Rain100L (Yang et al. 2017), Test2800 (Fu et al. 2017), and Test1200 (Zhang and Patel 2018). We employ the PSNR and SSIM (Wang et al. 2004) metrics to evaluate the restoration performance. For the image super-resolution and image deraining tasks, the metrics are computed on the Y channel in the YCbCr color space. For the image deblurring, PSNR and SSIM are evaluated in the RGB color space.
Implementation Details
For image super-resolution, the input is randomly cropped into $48\times 48$ patches and augmented by random horizontal and vertical flips and rotations. All the models are trained using ADAM optimizer (Kingma and Ba 2014) with $\beta_{1}=0.9$, $\beta_{2}=0.99$, and $\epsilon=10^{-8}$. The training batch size is set to 16 with a total of $2.5\times 10^{5}$ iterations. The initial learning rate is set to $10^{-4}$ and is decayed by a factor of 10 at every $10^{5}$ update. DCKD is implemented by PyTorch using 4 NVIDIA V100 GPUs. For other image restoration tasks, we strictly adhere to the original training configurations of each teacher model. The training configurations for other tasks are detailed in the Appendix.
Results and Comparison
Scale Method Set5 Set14 BSD100 Urban100
PSNR PSNR PSNR PSNR
$\times$2 Teacher 38.27 34.13 32.41 33.34
Scratch 38.13 33.78 32.26 32.63
Logits 38.17 33.83 32.29 32.67
FAKD 38.17 33.83 32.29 32.65
CrossKD 38.18 33.81 32.30 32.66
MiPKD 38.21 33.92 32.32 32.83
DCKD 38.25 34.01 32.35 32.92
$\times$3 Teacher 34.74 30.65 29.32 29.09
Scratch 34.61 30.45 29.18 28.59
Logits 34.61 30.47 29.21 28.62
FAKD 34.63 30.51 29.21 28.62
CrossKD 34.65 30.50 29.21 28.60
MiPKD 34.72 30.55 29.25 28.76
DCKD 34.74 30.60 29.29 28.87
$\times$4 Teacher 32.63 28.87 27.77 26.82
Scratch 32.38 28.69 27.63 26.36
Logits 32.45 28.76 27.67 26.49
FAKD 32.46 28.77 27.68 26.50
CrossKD 32.46 28.79 27.69 26.52
MiPKD 32.46 28.77 27.69 26.55
DCKD 32.56 28.82 27.73 26.69
$\mathrm{\textbf{DCKD}}^{*}$ 32.58 28.86 27.74 26.74
Table 2: Quantitative comparison of the benchmark datasets for the RCAN model. The best and second-best performances are highlighted in bold and underlined, respectively. DCKD* adds the optimized upper bound constraint.
Model Method GoPro #Params
PSNR/SSIM
MT-RNN - 31.15/0.9450 2.6M
DMPHN - 31.20/0.9400 21.7M
NAFNet Teacher 32.87/0.9606 17.1M
Scratch 31.17/0.9457 -
Logits 31.26/0.9464 -
DCKD 31.43/0.9487 2.7M
Restormer Teacher 32.92/0.9610 26.1M
Scratch 31.57/0.9497 -
Logits 31.61/0.9501 -
DCKD 31.78/0.9521 3.8M
Table 3: Quantitative comparison on the GoPro dataset for the deblurring task.
Figure 4: The image deblurring visual comparison of NAFNet on the GoPro dataset.
Image Super-Resolution
We compare our framework with the representative KD methods: train from scratch, Logits (Hinton, Vinyals, and Dean 2015), FAKD (He et al. 2020), CrossKD (Fang et al. 2023), and MiPKD (Li et al. 2024), on ×$\times 2$, ×$\times 3$, and ×$\times 4$ super-resolving scales.
The quantitative results for transformer-based backbone SwinIR and CNN-based backbone RCAN are presented in Tab. 1 and Tab. 2, respectively. Existing KD methods provide limited improvement for the student model, and in some cases, they even lead to worse performance compared to models trained without KD. For example, using FAKD for distillation on Urban100 results in poorer performance than training the model from scratch on SwinIR. DCKD is effective for both Transformer-based and CNN-based architectures, significantly outperforming existing KD methods by more than 0.1dB across all three scales on the Urban100 dataset. We deliberately used the most straightforward approach to demonstrate DCKD, showing that dynamic lower-bound constraints can yield strong results even without improving upper-bound constraints. To demonstrate that DCKD can be combined with methods that optimize upper-bound constraints, we incorporated DUKD (Zhang et al. 2023) into the DCKD framework to enhance upper-bound constraints, resulting in DCKD*. As we can see, the proposed DCKD can be combined with existing KD methods that optimize the upper bound, further significantly enhancing performance. DCKD* significantly outperforms the SOTA method MiPKD by 0.25dB on SwinIR and 0.19dB on RCAN at ×$\times 4$.
Figure 5: The ×$\times 4$ super-resolution visual comparison of SwinIR on the Urban100 dataset.
Figure 6: The ×$\times 4$ super-resolution visual comparison of RCAN on the Urban100 dataset.
Model Method Test100 Rain100H Rain100L Test2800 Test1200 #Params
PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM
MPRNet - 30.27/0.8970 30.41/0.8900 36.40/0.9650 33.64/0.9380 32.91/0.9160 20.1M
SPAIR - 30.35/0.9090 30.95/0.8920 36.93/0.9690 33.34/0.9360 33.04/0.9220 -
Restormer Teacher 32.02/0.9237 31.48/0.9054 39.08/0.9785 34.21/0.9449 33.22/0.9270 26.1M
Scratch 31.01/0.9122 30.51/0.8932 37.47/0.9714 33.78/0.9396 33.67/0.9295 -
Logits 31.04/0.9143 30.48/0.8915 37.17/0.9712 33.81/0.9399 33.78/0.9310 -
Our 31.08/0.9167 30.54/0.8969 38.02/0.9762 33.91/0.9411 33.95/0.9326 3.8M
Table 4: Quantitative comparison of the benchmark datasets for the deraining task.
Fig. 5 and Fig. 6 present challenging visual examples for Transformer and CNN backbones, respectively. Compared to existing KD methods, our approach enables the student model to better capture structural textures, such as more accurately reconstructing sidewalk lines and building structures. More visual comparisons for various examples and models are presented in the Appendix.
Image Deblurring
Tab. 3 provides the quantitative comparison on the GoPro dataset (Nah, Hyun Kim, and Mu Lee 2017b). Our method demonstrates consistent effectiveness on both the CNN-based NAFNet (Chen et al. 2022a) and the Transformer-based Restormer (Zamir et al. 2022). Compared to the Logits KD, our DCKD achieves 0.17dB performance improvement on different backbones. Moreover, with comparable parameters, the student model of NAFNet significantly outperforms MT-RNN (Park et al. 2020) by 0.28dB. Fig. 4 illustrates the deblurring visualization results. DCKD restores the clearest window outlines, significantly enhancing the deblurring capability of the student model.
Figure 7: The image deraining visual comparison of Restormer on the Rain100L dataset.
Image Deraining
Tab. 4 shows the performance of different methods on several benchmark datasets for image draining. With only 18.9% of MPRNet’s parameters, DCKD significantly surpasses it by 1.8dB in PSNR on the Rain100L dataset. Additionally, compared to other distillation methods, DCKD outperforms Logits KD by 0.55dB on the Rain100L. Fig. 7 presents a visual comparison of derained images. DCKD further enhances the student’s ability to remove rain streaks compared to the logit distillation method.
DCR DMM Set14 Urban100
PSNR/SSIM PSNR/SSIM
33.83/0.9197 32.67/0.9329
33.98/0.9208 32.83/0.9346
33.92/0.9205 32.81/0.9343
34.01/0.9213 32.92/0.9351
Table 5: Ablation study on components of our framework.
Degradation Type Set14 Urban100
PSNR/SSIM PSNR/SSIM
Random Blur 34.01/0.9212 32.89/0.9352
Random Resize 33.99/0.9212 32.90/0.9354
Random Noise 34.01/0.9213 32.92/0.9351
Random Mix 34.00/0.9212 32.87/0.9353
Table 6: Ablation study on the type of degradation.
$\lambda_{dcl}$ 0.01 0.1 1.0
PSNR/SSIM 32.87/0.9346 32.92/0.9351 32.81/0.9350
$\lambda_{ce}$ 0.0001 0.001 0.01
PSNR/SSIM 32.84/0.9345 32.92/0.9351 32.82/0.9341
Table 7: Ablation study on the balancing weights of our framework.
Figure 8: Ablation studies on the number of negative samples and the initial update step.
Ablation Study
For ablation experiments, we train DCKD on RCAN (Zhang et al. 2018) with the ×$\times 2$ scale for the SR task. We then validate the results on the Set14 (Zeyde, Elad, and Protter 2012) and Urban100 (Huang, Singh, and Ahuja 2015).
Components of the Proposed Framework
As Tab. 5, we first conduct an ablation study on the two main modules of DCKD. The results show that our DCR and DMM outperform the baseline by 0.16dB and 0.14dB on the Urban100, respectively. This demonstrates the effectiveness of the proposed modules. Additionally, the combination of DCR and DMM further enhances model performance, achieving improvements of 0.18dB on the Set14 and 0.25dB on the Urban100 compared to the baseline.
Type of Random Degradation
We conduct ablation studies on the type of degradation module in the Dynamic Negative Sample Generator (DNSG), including random blurring, random resize, random noise, and mixed degradation. As shown in Tab. 6, random noise achieves the best results, while random blurring and random resize produce a similar performance. The mixed degradation results in lower-quality negative samples, which weakens the lower bound constraint and leads to the worst performance. Consequently, we adopted random noise degradation in DCKD.
Impact of Balancing Weights
We investigate the impact of the balancing coefficients $\lambda_{dcl}$ and $\lambda_{ce}$ in Equ. 10, as shown in Tab. 7. We find that excessively large or small values for these coefficients negatively affect the outcomes. The experiments indicate that the model achieves optimal results when $\lambda_{dcl}$ is set to 0.1 and $\lambda_{ce}$ to 0.001. Given the broad applicability of our method to various image restoration tasks, we adopt $\lambda_{dcl}=0.1$ and $\lambda_{ce}=0.001$ as the default settings across different tasks.
Impact of the Number of Negative Samples
The impact of the number of negative samples is reported in Fig. 8 (a). The results indicate that as the number of negative samples increases, the performance improves consistently. However, when the number of negative samples exceeds 5, the performance gains diminish while significantly increasing training time and memory costs. Therefore, we set the number of negative samples to 5.
Impact of Initial Update Step
In Fig. 8 (b), we investigate the impact of the initial update step for the history model within the dynamic negative sample generator. The experimental results show that when using a smaller step to update the historical model, the quality of the negative samples becomes very close to, or even surpasses, that of the anchor points, leading to instability in the solution space and a decline in performance. Conversely, when using a larger step, the negative sample quality deteriorates, weakening the lower bound constraint. When the initial update step is set to 1000 achieve the best performance.
Conclusion
In this work, we propose a dynamic contrastive knowledge distillation framework for image restoration, named DCKD, which consists of Dynamic Contrastive Regularization (DCR) and Distribution Mapping Module (DMM). Most previous knowledge distillation methods utilize a fixed solution space, causing the lower bound constraints to weaken gradually during training. DCR constructs a dynamic solution space based on the student’s learning state to enhance lower-bound constraints. DMM is the first time introducing pixel-level category information for knowledge distillation in low-level vision tasks. Experiments on image super-resolution, image deraining, and image deblurring tasks validate that the proposed DCKD achieves state-of-the-art results on various benchmark datasets both quantitatively and visually.